OpenStudy (anonymous):

8!/3! 72 7360 120 6720

OpenStudy (bahrom7893):

8!/3! = 8*7*6*5*4*3*2*1/(3*2*1)

OpenStudy (bahrom7893):

= 8*7*6*5*4 = 56*6*20 = 56 * 120 <-whatever that is

OpenStudy (anonymous):

thanks!